# Know Your Needs Better: Towards Structured Understanding of Marketer Demands with Analogical Reasoning Augmented LLMs

### v1

@@ Abstract.
In this paper, we explore a new way for user targeting, where non-expert marketers could select their target users solely given demands in natural language form. The key to this issue is how to transform natural languages into practical structured logical languages, i.e., the structured understanding of marketer demands. Considering the impressive natural language processing ability of large language models (LLMs), we try to leverage LLMs to solve this issue. Past research indicates that the reasoning ability of LLMs can be effectively enhanced through chain-of-thought (CoT) prompting. But existing methods still have some limitations: (1) Previous methods either use simple “Let’s think step by step” spells or provide fixed examples in demonstrations without considering compatibility between prompts and questions, making LLMs ineffective in some complex reasoning tasks such as structured language transformation. (2) Previous methods are often implemented in closed-source models or excessively large models, which is not suitable in industrial practical scenarios. Based on these, we propose ARALLM (i.e., Analogical Reasoning Augmented Large Language Models) consisting of two modules: Analogical Reasoning based Prompting and Reasoning-Augmented Multi-Task Model Distillation. To be specific, we first introduce a structured language SELL that is practical in marketing scenarios and construct a reasoning library consisting of several high-quality reasoning examples of translating natural language to SELL (NL2SELL). Then, we adopt a retrieval-based method to conduct analogical reasoning with the help of the reasoning library. The experimental results show that prompting with analogical reasoning achieves better performance than the ordinary prompting method in NL2SELL tasks. Beyond that, we distill knowledge from super LLMs (ChatGPT) based on analogical reasoning and construct a NL2SELL dataset to fine-tune smaller student LLMs in a multi-task training paradigm, enabling the models to be easily deployed in practical user targeting environments.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing
@@ 1. Introduction
Unfortunately, both of the above prior studies have primarily emphasized the intricate architectures for effectively and efficiently gathering targeted users, while commonly ignoring the natural and significant gap between marketers’ demand and the capability of current models. In particular, current user targeting pipelines mainly force the marketers to decompose their demands into multiple tags/entities, e.g., tags such as “Preference”, “Resident City” and “Age” are characterized in Figure 1. The process incurs several weaknesses: i) unfriendly to marketers; ii) time-consuming for manual deconstruction of demands; and iii) unpromising performance due to partial tags/entities.
@@
Therefore, in this paper, we take the initial stride towards enhancing current user targeting systems by probing into the structured understanding of marketers’ demands in an automatic way, enabling the system to know your needs better solely based on natural language inputs. As shown in Figure 1 (c), the marketers only need to input their native demand like “Young people in Hangzhou who enjoy milk tea or white-collar workers in first-tier cities who enjoy listening to audiobooks” and our new approach can instantly provide the user targeting cards through a novelty structured understanding of the demand. After verifying the cards, they can export the results with just one click and the system will automatically run the existing targeted user aggregation modules. The whole process only takes 3-4 minutes, which is 10 times faster than the former approach. The core of this new approach lies in how to transform natural languages into structured forms, or so-called structured understanding of marketer demands.
@@
To address this issue, we first develop a Structured and Editable Logical Language called SELL, which forms the basis of the user targeting cards shown in Figure 1(c). The main concern of our work is transforming the Natural Language into SELL (NL2SELL). For example, the natural language “Young people in Hangzhou who enjoy milk tea or white-collar workers in first-tier cities who enjoy listening to audiobook” can be translated into SELL as “((Resident City#Belongs To# Hangzhou) AND (Age#Between#18,35) AND (Preference#Belongs To#Milk Tea)) OR ((City Level#Belongs To#First-tier) AND (Days of listening to audiobooks#Greater Than#3) AND (Career#Belongs To#White-collar))”. Previous research  (Liu et al., 2023; Gu et al., 2023; Yang et al., 2023b; Lu et al., 2022) has highlighted the remarkable abilities of Large Language Models (LLMs) in language translation tasks. LLMs are renowned for their robust natural language processing and zero-shot capabilities, which render them possibly effective in NL2SELL. Consequently, we opt to employ LLMs to tackle this task.
@@
However, there are still some challenges in using LLMs in our scenario: (1) The challenge of reasoning accuracy. There is still room for improvement in reasoning accuracy for LLMs in complex problems, such as structured understanding of natural language and translation issues in our scenario. Recently, it (Wei et al., 2022; Kojima et al., 2022) has been pointed out that prompting LLMs with chain-of-thought (CoT) could enhance the reasoning ability of LLMs. These methods either provide fixed reasoning examples as demonstrations through few-shot learning or tell the model “Let’s think step by step” for zero-shot learning. Though effective, they ignore the compatibility between prompts and specific questions. In some challenging reasoning scenarios, zero-shot learning or having few related reasoning examples in few-shot learning may not effectively stimulate the LLMs’ reasoning ability (Rubin et al., 2021). Ideally, we can customize a reasoning process for each question to induce the model to reason better, but it is time-consuming and laborious. (2) The challenge of reasoning speed and resource consumption. Although the reasoning ability of super-large language models can be developed through prompt engineering, it is unacceptable to deploy excessively huge LLMs or API-based LLMs in practical situations. For extremely huge LLMs, excessive parameters will consume a lot of resources, and the reasoning speed is also slower. For API-based LLMs such as ChatGPT, both the security of private data and the cost of API calls are thorny issues. Thus, smaller, free, and white-box LLMs are preferred.
@@
To address these challenges, we propose a framework called Analogic Reasoning Augmented Large Language Model (ARALLM) to tackle the NL2SELL task, which consists of two modules: Analogical Reasoning based Prompting and Reasoning-Augmented Multi-Task Model Distillation. (1) Analogical Reasoning based Prompting. Creating a CoT for each new targeting demand is impractical and time-consuming (considering the need for deployment). Thus, we propose Analogical Reasoning based Prompting. The fundamental concept of it is that when two targeting demands are similar, they might follow similar reasoning steps and have similar conditional expression structures. Therefore, we construct a compact yet reliable reasoning library to serve as a reference for numerous demands that have unknown reasoning steps. This approach helps reduce the effort of manually writing reasoning steps. Additionally, we can offer assistance for any unfamiliar demands by employing analogical reasoning with retrieval. (2) Reasoning-Augmented Multi-Task Model Distillation. Since the white-box and smaller model is preferred, we first distill knowledge from teacher LLMs(ChatGPT) to construct a NL2SELL dataset with over 10,000 pieces of training samples and then propose a novel multi-task training method with reasoning augmentation.
@@
We summarize our contributions as follows: (1) Aiming to automatically understand marketers’ demands in a structured manner, we propose a brand-new language SELL and NL2SELL task. (2) We propose an innovative framework called ARALLM, which is composed of the Analogical Reasoning based Prompting module and the Reasoning-Augmented Multi-Task Distillation module, to address the NL2SELL task. (3) We demonstrate the superiority of the proposed ARALLM framework through extensive experiments on real-world datasets.
2. Design of SELL
In this section, we will provide a detailed explanation of the design and the usage of the SELL. The SELL is mainly composed of four elements:
Keys are series of tags that describe the features of the user, such as Age, Gender, Monthly Income, Pet Owning, and so on. In practical marketing scenarios, the database contains tens of thousands of tags that describe user profiles. Selecting the appropriate keys from this vast array of tags is a challenging task.
Values are the fillings of the corresponding keys. They can be generally divided into three types: numerical type, string type, and boolean type. For instance, the value corresponding to the key Monthly Income is of numerical type, because they are continuous and non-enumerable. The value of Gender is of string type, with limited options: Male or Female. Boolean type values have only two states, i.e., True or False, such as Pet Owning.
Operators represent the relationship between keys and values. For numerical values, seven types of operators can be used to connect them with corresponding keys: Equal To, Greater Than, Less Than, Not Equal To, Not Greater Than, Not Less Than, and Between. For values of string type and boolean type, there are two types of operators: Belongs To and Not Belongs To. Keys, operators, and values can form a basic conditional expression in SELL, formatted as “(key#operator#value)”, such as “(Pet Owning#Belongs To#True)”. A conditional expression represents a cluster of target users, for example, in the above example, the target users are pet owners.
Intersection and Union symbols are used to take the intersection or union of multiple conditional expressions. In practical marketing scenarios, the features of the target users are often complex and difficult to describe using a single conditional expression. Therefore, it is necessary to combine the target users through intersection or union operations. In SELL, we use the symbols AND and OR to denote the intersection and union operations, respectively. For example, “((Resident City#Belongs To#Hangzhou) OR (Resident City#Belongs To#Shanghai)) AND (Pet Owning#Belongs To#True)” describes the pet owners who live in Hangzhou or Shanghai.
Compared to the popular database languages SQL, the advantages of SELL are: (1) SELL is simpler in expression and has much fewer syntax symbols than SQL, making it easier for non-programming experts to understand. (2) Since SELL can be transformed into a clear tree structure, marketers can easily modify them even if the generated results have flaws.
@@ 3. Method
In this work, we focus on how to convert a sentence of marketer’s demand $d$ into a standard SELL expression, while how to parse and execute the SELL and return the target users from the database will be handled by the internal system, which does not constitute the core of this article. Inspired by the analogical reasoning (Yao et al., 2023; Turney, 2008; Gentner, 1983) in the classic artificial intelligence field and CoT prompting methods in LLMs, we first propose a novel prompting method enhanced by analogical reasoning and then adopt a reasoning-augmented multi-task training strategy to fine-tune smaller LLMs.
@@ 3.1. Prompting With Analogical Reasoning
The core idea of analogical reasoning in the NL2SELL task is that if two demands are similar, they may share similar reasoning steps, as well as similar conditional expression structures. Thus, we construct a small but high-quality reasoning library to provide references for numerous demands with unknown reasoning steps, minimizing the cost of manually writing reasoning steps for each of them.
@@ 3.1.1. Construction of Reasoning Library
To construct a reasoning library, we first randomly select 50 marketer demands from real-world scenarios and write their corresponding answers based on our expert knowledge to form Q&A pairs =$\mathcal{R}=\{(d_{e},s_{e})\}_{e=1}^{50}$, where $d_{e}$ is the demand, $s_{e}$ is the corresponding SELL expressions. Thanks to the simplicity of the SELL itself, this processing can be completed by two experts within an hour. Then we use a small amount of manual effort to write corresponding reasoning details for 5 different ($d_{e}$,$s_{e}$) $\in \mathcal{R}$ to construct the seed examples in the format $(d_{e},r_{e},s_{e})$, where $r_{e}$ represent the reasoning details. We generally summarize the procedure of solving the NL2SELL problem as four steps:
@@
For the remaining data in $\mathcal{R}$, we utilize the OpenAI gpt-3.5-turbo API [^1] https://platform.openai.com/docs/api-reference/chat to help us fill the corresponding reasoning steps, and our seed examples serve as demonstrations in the prompt. After simple verification, we harvest a batch of solid reasoning details to form a reasoning library =$\mathcal{R}=\{(d_{e},r_{e},s_{e})\}_{e=1}^{50}$.
@@ 3.1.2. Retrieval-Based Analogical Prompt Construction
where E is the embedding model which we use BGE [^2] https://huggingface.co/BAAI/bge-large-zh(Xiao et al., 2023; Zhang et al., 2023) off the shelf as it is state-of-the-art in Chinese text retrieval field. The top $k$ demands in $\mathcal{R}$ that get the highest similarity score when compared with $d_{p}$ will be fetched as analogical examples in this stage, along with their corresponding $r$ and $s$.
@@
Compared to the ordinary CoT method, our approach exhibits several main differences:
First, the NL2SELL task is more challenging. Unlike solving mathematical problems or common sense reasoning tasks, the training data for LLMs don’t inherently include the syntax of SELL. This highlights the importance of appropriate reasoning steps.
Second, we place a greater emphasis on the compatibility between prompts and questions. We believe that using analogical reasoning based prompting method can make LLMs more adaptable to specific problems, rather than relying on fixed few-shot examples, which could potentially limit the model’s performance.
@@ 3.2. Reasoning-Augmented Multi-Task Model Distillation
To train a private small model for deployment, we first distill knowledge from teacher LLM (ChatGPT) to construct a NL2SELL dataset with over 10,000 pieces of training samples and then propose a multi-task training method with reasoning augmentation.

### v3

@@ Abstract.
In this paper, we explore a new way for user targeting, where non-expert marketers could select their target users solely given demands in natural language form. The key to this issue is how to transform natural languages into practical structured logical languages, i.e., the structured understanding of marketer demands. In practical scenarios, the demands of non-expert marketers are often abstract and diverse. Considering the impressive natural language processing ability of large language models (LLMs), we try to leverage LLMs to solve this issue. To stimulate the LLMs’ reasoning ability, the chain-of-thought (CoT) prompting method is widely used, but existing methods still have some limitations in our scenario: (1) Previous methods either use simple “Let’s think step by step” spells or provide fixed examples in demonstrations without considering compatibility between prompts and concrete questions, making LLMs ineffective when the marketers’ demands are abstract and diverse. (2) Previous methods are often implemented in closed-source models or excessively large models, which is not suitable in industrial practical scenarios. Based on these, we propose ARALLM (i.e., Analogical Reasoning Augmented Large Language Models) consisting of two modules: Analogical Reasoning based Prompting and Reasoning-Augmented Multi-Task Model Distillation. To be specific, we first construct a reasoning library consisting of several high-quality and topic-rich reasoning examples. Then, we adopt a retrieval-based method to conduct analogical reasoning with the help of the reasoning library. The experimental results show that this prompting strategy achieves better performance than the ordinary prompting method. Beyond that, we distill knowledge from super LLMs (GPT-3.5) to fine-tune smaller student LLMs in a multi-task training paradigm, enabling the models to be easily deployed in practical environments. Part of our data and code can be found at https://github.com/alipay/Analogic-Reasoning-Augmented-Large-Language-Model.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671583isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Natural language processing
@@ 1. Introduction
Unfortunately, both of the above prior studies have primarily emphasized the intricate architectures for effectively and efficiently gathering targeted users, while commonly ignoring the natural and significant gap between marketers’ demand and the capability of current models. In particular, current user targeting pipelines mainly force the marketers to decompose their demands into multiple tags/entities, e.g., “Preference”, “Resident City” and “User Age Group” which are characterized in Figure 1. The process incurs several weaknesses: i) unfriendly to marketers; ii) time-consuming for manual deconstruction of demands; and iii) unpromising performance due to partial tags/entities.
@@
Therefore, in this paper, we take the initial stride towards enhancing current user targeting systems by probing into the structured understanding of marketers’ demands in an automatic way, enabling the system to know your needs better solely based on natural language inputs. As shown in Figure 1 (c), the marketers only need to input their native demands like “Young people in City A who enjoy milk tea or white-collar workers in first-tier cities who enjoy listening to audiobooks” and our new approach can instantly conduct structured decomposition of marketers’ demands and thus provide the user targeting cards. After verifying the cards, they can export the target users with just one click and the system will automatically run the existing targeted user aggregation modules.
@@
The core of this new approach lies in how to transform natural languages into structured forms, or so-called structured understanding of marketer demands. To address this issue, we first draw upon existing logical and programming languages to design a logical expression that offers both readability and practical applicability for the structured representation of marketers’ demands. We name it as SELL, i.e., Structured and Editable Logical Language, which mainly consists of Keys (e.g., Resident City), Values (e.g., City A), Operators (e.g., Belongs To) and Intersection/Union symbols (i.e., AND, OR). It forms the basis of the user targeting cards shown in Figure 1(c). The main concern of our work is transforming the Natural Language into SELL (NL2SELL). For example, the natural language “Young people in City A who enjoy milk tea or white-collar workers in first-tier cities who enjoy listening to audiobook” can be translated into SELL as “((Resident City#Belongs To# City A) AND (User Age Group#Between#18,35) AND (Preference#Belongs To#Milk Tea)) OR ((City Level#Belongs To#First-tier) AND (Days of listening to audiobooks#Greater Than#3) AND (Career#Belongs To#White-collar))”.
@@
Previous research  (Liu et al., 2023; Gu et al., 2023; Yang et al., 2023b; Lu et al., 2022) has highlighted the remarkable abilities of Large Language Models (LLMs) in language translation tasks. LLMs are renowned for their robust natural language processing and zero-shot capabilities, which render them possibly effective in NL2SELL. Consequently, we opt to employ LLMs to tackle this task. However, there are still some challenges in using LLMs in our scenario: (1) The challenge of reasoning accuracy. In practical user targeting scenarios, understanding of marketer’s demands could be very challenging. For example, a marketer might want to market products related to education, targeting parents of middle school students who are focused on education. Therefore, the demand she inputs into the system is a simple sentence “Parents of middle school students”, and we need to use existing tags to convert this demand into a structured expression, such as “(Marital Status#Belongs#True) AND (User Child Age#Between#12,15) AND (Preference#Belongs To#Education)”. However, the key “Marital Status, User Child Age, and Preferences” do not directly appear in the marketer’s demand, which requires the LLMs to possess a high level of language comprehension and reasoning abilities. Recently, it (Wei et al., 2022; Kojima et al., 2022) has been pointed out that prompting LLMs with chain-of-thought (CoT) could enhance the reasoning ability of LLMs. These methods either provide fixed reasoning examples as demonstrations through few-shot learning or tell the model “Let’s think step by step” for zero-shot learning. Though effective, they ignore the compatibility between prompts and specific questions. Considering the diversity of marketers’ demands themselves, zero-shot learning or having fixed reasoning examples in few-shot learning may not effectively stimulate the LLMs’ reasoning ability (Rubin et al., 2021). Ideally, we can customize a reasoning process for each question to induce the model to reason better, but it is time-consuming and laborious. (2) The challenge of reasoning speed and resource consumption. Although the reasoning ability of super-large language models can be developed through prompt engineering, it is unacceptable to deploy excessively huge LLMs or API-based LLMs in practical situations. For extremely huge LLMs, excessive parameters will consume a lot of resources, and the reasoning speed is also slower. For API-based LLMs such as GPT-3.5, both the security of private data and the cost of API calls are thorny issues. Our model needs to be deployed in real ToC (To Customer) scenarios, where it will generate at least hundreds of tasks daily. Therefore, there is an urgent requirement to increase the inference speed, reduce the cost of inference, and ensure data security. Thus, smaller, free, and white-box LLMs are preferred.
@@
To address these challenges, we propose a framework called Analogic Reasoning Augmented Large Language Model (ARALLM) to tackle the NL2SELL task, which consists of two modules: Analogical Reasoning based Prompting and Reasoning-Augmented Multi-Task Model Distillation. (1) Analogical Reasoning based Prompting. Creating a CoT for each new targeting demand is impractical and time-consuming (considering the need for deployment). Thus, we propose Analogical Reasoning based Prompting. The fundamental concept of it is that when two targeting demands are similar, they might follow similar reasoning steps and have similar conditional expression structures. Therefore, we construct a compact yet reliable reasoning library to serve as a reference for numerous demands that have unknown reasoning steps. This approach helps reduce the effort of manually writing reasoning steps for every demand. Additionally, we can offer assistance for any unfamiliar demands by employing analogical reasoning with retrieval. (2) Reasoning-Augmented Multi-Task Model Distillation. Since the white-box and smaller model is preferred, we distill knowledge from teacher LLMs (GPT-3.5) to construct an NL2SELL dataset with over 10,000 pieces of training samples and then propose a novel multi-task training method with reasoning augmentation to improve the reasoning ability of student LLMs.
@@
We summarize our contributions as follows: (1) We explore a new way for marketer-friendly user targeting, which focuses on the structured understanding of marketers’ raw demands. (2) We propose an innovative framework called ARALLM, which is composed of the Analogical Reasoning based Prompting module and the Reasoning-Augmented Multi-Task Distillation module, to address the above task. (3) We demonstrate the superiority of the proposed ARALLM framework through extensive experiments on real-world datasets and make this LLMs-based framework operate in real industry scenarios.
@@ 2. Method
In this work, we focus on how to convert a sentence of marketer’s demand $d$ into a standard structured expression SELL, while how to parse and execute the SELL and return the target users from the database will be handled by the internal system, which does not constitute the core of this article. Inspired by the analogical reasoning (Yao et al., 2023; Turney, 2008; Gentner, 1983) in the classic artificial intelligence field and CoT prompting methods in LLMs, we first propose a novel prompting method enhanced by analogical reasoning and then adopt a reasoning-augmented multi-task training strategy to fine-tune smaller LLMs.
2.1. Design of SELL
We first explain the design of SELL to help readers understand our task better. The SELL is mainly composed of four elements: Keys, Values, Operators, and Intersection and Union symbols:
(1) Keys are a series of tags that describe the features of the user, such as Gender, Monthly Income, Pet Owning, and so on.
(2) Values are the fillings of the corresponding keys. They can be generally divided into three types: numerical type, string type, and boolean type.
(3) Operators represent the relationship between keys and values. For numerical values, seven types of operators can be used to connect them with corresponding keys: Equal To, Greater Than, Less Than, Not Equal To, Not Greater Than, Not Less Than, and Between. For values of string type and boolean type, there are two types of operators: Belongs To and Not Belongs To. Keys, operators, and values can form a basic conditional expression in SELL, formatted as “(key#operator#value)”, such as “(User Marital Status#Belongs#True)”. A conditional expression represents a cluster of target users.
(4) Intersection and Union symbols are used to take the intersection or union of multiple conditional expressions (targeted users). We use the symbols AND and OR to denote the intersection and union operations, respectively.
Compared to the popular database languages SQL and logic language FOL, the advantages of SELL are: (1) SELL is simpler in expression and has much fewer syntax symbols, making it easier for non-programming experts to understand. (2) Since SELL can be transformed into a clear tree structure, marketers can easily modify them even if the generated results have flaws. Due to space limitations, we recommend readers to read more details about SELL in the supplementary materials.
@@ 2.2. Prompting With Analogical Reasoning
The core idea of analogical reasoning in the NL2SELL task is that if two demands are similar, they may share similar reasoning steps, as well as similar logical structures. Thus, we construct a small but high-quality reasoning library to provide references for numerous demands with unknown reasoning steps, minimizing the cost of manually writing reasoning steps for each of them.
@@ 2.2.1. Construction of Reasoning Library
To construct a reasoning library, we first randomly collect $\mathcal{N}_{RL}$ marketer demands from real-world scenarios and write their corresponding answers based on our expert knowledge to form Q&A pairs =$\mathcal{R}=\{(d_{e},s_{e})\}_{e=1}^{\mathcal{N}_{RL}}$, where $d_{e}$ is the demand, $s_{e}$ is the corresponding SELL expressions. The collected marketer’s demands encompass services that are popular on Alipay, including education, technology, reading, travel, financial management, insurance, government affairs, and public welfare, thus ensuring that the reasoning library constructed based on this can cover as many demand topics as possible. Then we use a small amount of manual effort to write corresponding reasoning details for 10% random selected Q&A pairs $\in \mathcal{R}$ to construct the seed examples in the format $(d_{e},r_{e},s_{e})$, where $r_{e}$ represent the reasoning details. We generally summarize the procedure of solving the NL2SELL problem as four steps:
@@
For the remaining data in $\mathcal{R}$, we utilize the OpenAI gpt-3.5-turbo-1106 (GPT-3.5) API to help us fill the corresponding reasoning steps, and our seed examples serve as demonstrations in the prompt. Since we have standardized the overall framework of reasoning, it is not difficult for GPT-3.5 to complete this part of the reasoning steps, cases are shown in supplemental materials. After being verified by experts. we harvest a batch of solid reasoning details to form a reasoning library =$\mathcal{R}=\{(d_{e},r_{e},s_{e})\}_{e=1}^{\mathcal{N}_{RL}}$.
@@ 2.2.2. Retrieval-Based Analogical Prompt Construction
where E is the embedding model in which we use BGE-large-zh (Xiao et al., 2023; Zhang et al., 2023) off the shelf as it is well-performed in the Chinese text retrieval field. The top $k$ demands in $\mathcal{R}$ that get the highest similarity score when compared with $d_{p}$ will be fetched as analogical examples in this stage, along with their corresponding $r$ and $s$.
@@
Compared to the ordinary CoT method, our approach exhibits several main differences: (1) The NL2SELL task is more challenging. Unlike solving mathematical problems or common sense reasoning tasks, the training data for LLMs don’t inherently include the syntax of SELL. This highlights the importance of appropriate reasoning steps. (2) We place a greater emphasis on the compatibility between prompts and questions. We believe that using an analogical reasoning based prompting method can make LLMs more adaptable to specific problems, rather than relying on fixed few-shot examples, which could potentially limit the model’s performance.
@@ 2.3. Reasoning-Augmented Multi-Task Model Distillation
To train a private small model for deployment, we first distill knowledge from teacher LLM (GPT-3.5) to construct an NL2SELL dataset with over 10,000 pieces of training samples and then propose a multi-task training method with reasoning augmentation.